A positive number is 5 time another number. If 21 is added to both the numbers, then one of the new numbers become twice of other new numbers. Find the original numbers.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the smaller number be x
And another number = 5x

If 21 is added to both the numbers then as per the given condition;

5x+21 = 2×(x+21)

5x + 21 = 2x + 42

5x – 2x = 42 – 21

3x =. 21

x = 21/3= 7

Therefore, the numbers are = 35 and 7

Thanks